Problem 1 (20 points) Given the following non-linear autonomous system, || x = 2cy || y = 9-+ y2 : a) What are the equili
Answer #1

x = 2x4 y=9-x²+42 1. X 2=0, y = -9 w (a) The equilibrium points are, x=0, y=0 = 22y =o > Boys xa 2= 0 on 4 = 0 5 :: = 0, T重重重重重重重重重重重 (6 0A|| | so » 13 so A = + Fi? (-3,0) is stable limit cycle. 影 - 278 = =-x+y^ > 24 影 - -x+y | let, i=str>] = 最 能